OpenStudy (anonymous):

find the value of 59^2

OpenStudy (solomonzelman):

3481 google calculator

OpenStudy (anonymous):

4 59 *59 ---- 531 295 ------- 3481

OpenStudy (aum):

\(59^2 = (60-1)^2 = 60^2 -2*60*1 + 1 = 3600 - 120 + 1 = 3481 \)

OpenStudy (solomonzelman):

another approach 59² is like 60² but -(60+59)

OpenStudy (anonymous):

Evaluate if x=12 and y=10 : 3xy-2x+7y

OpenStudy (solomonzelman):

And so is 79²=80²-80-79

OpenStudy (aum):

Replace x by 12 and y by 10 in 3xy - 2x + 7y and evaluate the expression.